# TrendFact: A Benchmark for Explainable Hotspot Perception in Fact-Checking with Natural Language Explanation

### v3

@@ Appendix C Fact-Checking Claim Rewriting Factor
As shown in Table 8, six critical factors are proposed to systematically transform social media headlines into verifiable claims: (1) Temporal Anchoring converts vague temporal expressions into specific time references; (2) Data Granularity decomposes aggregated data into measurable units; (3) Ambiguity Resolution replaces probabilistic terms with deterministic statements; (4) Comparative Standard introduces quantifiable benchmarks; (5) Domain Knowledge integrates industry-specific parameters; and (6) Source Implication embeds provenance cues. These factors collectively enhance claim verifiability while preserving semantic coherence for automated processing. Concrete rewriting examples are illustrated in Figure .
Figure 6: Annotate Evaluation Criteria.
@@ Appendix E Evidence Library Construction Process
Figure 7 illustrates the detailed evidence library construction process for TrendFact. First, we employ GPT-4o to rewrite existing claims, with Figure 8 showcasing the specific rewriting prompt. Specifically, we transform original statements into challenging variants by maintaining core fact relevance (relevance score >0.6) and introducing at least three types of semantic alterations: subject displacement (e.g., institution → individual), causal inversion (e.g., "A leads to B" → "B triggers A"), degree conversion (e.g., "significant growth" → "slight fluctuation"), spatiotemporal shift (across years/regions), and quantification method change (absolute value → percentage), allowing for reasonable logical leaps. The rewritten claims are then processed through Bing web retrieval, retaining the top 10 search results and extracting their textual content. Subsequently, we merge and deduplicate the newly collected evidence with pre-existing gold evidence, followed by publication date crawling for each webpage to finalize the evidence library.
Appendix F Details of the ECS and HCPI
Table 9 provides a detailed definition of ECS. The descriptions from top to bottom are as follows: Dual Discrepancy: Both authenticity label misjudgment and fully inconsistent explanatory content. Label Error with Content Consistency: Incorrect authenticity labeling despite congruent explanatory material. Accurate Labeling with Explanatory Divergence: Correct authenticity identification accompanied by conflicting interpretation content. Partial Content Alignment: Proper authenticity classification with only partial consistency in explanatory elements. Full Verification Compliance: Complete congruence between correctly identified authenticity labels and their corresponding explanatory content.
The formulaic definition of the influence score in the HCPI metric is as follows:
$\text{Inf}(i)=r_{i}\cdot\left(\begin{aligned} &\alpha\cdot\log(1+v_{i})+\beta\cdot\log(1+d_{i})\\ &+\kappa\cdot\log(1+e_{i})+\lambda\cdot\log(1+p_{i})\end{aligned}\right)$ (2)
The influence score $\text{Inf}(i)$ of the i-th sample is calculated first, where $v_{i}$, $d_{i}$, $e_{i}$ and $p_{i}$ represent the views, discussions, engagements, and posts of the i-th sample, respectively. $\alpha$, $\beta$, $\kappa$ and $\lambda$ are the corresponding weights that sum up to 1. $r_{i}$ represents the risk index determined by GPT-4o. After obtaining the influence score, we calculate the HCPI metric as follows:
$\text{HCPI}=\frac{\begin{aligned} &\sum_{\mathclap{i\in\text{SUP}}}\,\begin{cases}\text{Inf}(i)\cdot\mathrm{ECS}(i),&\hat{y}_{i}=\text{SUP}\\ -2\cdot\text{Inf}(i),&\hat{y}_{i}=\text{REF}\end{cases}\\ &+\sum_{\mathclap{j\in\text{REF}}}\,\text{Inf}(j)\cdot\mathrm{ECS}(j)\cdot\delta(\hat{y}_{j}=\text{REF})\\ &+\sum_{\mathclap{k\in\text{NEI}}}\,\begin{cases}-\text{Inf}(k),&\hat{y}_{k}=\text{REF}\\ \text{Inf}(k)\cdot\mathrm{ECS}(k),&\hat{y}_{k}=\text{NEI}\end{cases}\end{aligned}}{\sum\limits_{m\in\text{All Claims}}\text{Inf}(m)}$ (3)
where SUP, REF, and NEI represent the three categories of fact-checking labels: Support, Refute, and Not Enough Information, respectively. ECS(i) denotes the Explanation Consistency Score for i-th sample, and $m$ is the total number of samples.
Label Verification Accuracy Explanation Consistency Score Normalized Score
Misjudgment Full Discrepancy 1 0.2
Misjudgment Consistency 2 0.4
Correct Judgment Content Divergence 3 0.6
Correct Judgment Partial Consistency 4 0.8
Correct Judgment Full Consistency 5 1.0
Table 9: Explanation Consistency Score.
@@ Appendix G Prompt of FactISR
As shown in Figure 9, the prompt details the following aspects: 1.EDA: Detailed Process of Dynamic Evidence Addition 2.ET: (1)The relevance between the claim and evidence. (2)The consistency between key information and evidence. (3)The conflict among the claim, evidence, label, and explanation.
@@
Appendix H Example of Rethinking via Reward Decoding
Figure 11 illustrates an example of reflection via reward decoding. Without reward decoding, the model directly outputs a conclusion of no conflict and prematurely ends the reasoning process. Our reward decoding encourages the model to reassess its previous judgment, leading to a reconsideration that ultimately results in the correct outcome.
Figure 11: An Example of Reflection via Reward Decoding.
@@ Appendix I Experiments Under Gold Evidence Conditions
The reward vector $\boldsymbol{R}$ and the decay factor $\gamma$ are set to 20 and 0.1, respectively. The maximum input length is set to 16k, while the maximum output lengths for LLMs and RLMs are set to 300 and 5k, respectively. The maximum length of retrieved results is truncated at 3k. The maximum number of retrievals is set to 3, and to ensure fair comparison, the maximum number of dynamically added evidences by our DEA is also limited to the same. All inference experiments utilized greedy search as the strategy. In this paper, for the HCPI metric, the values of $\alpha$, $\beta$, $\kappa$ and $\lambda$ used to calculate the influence score are set to 0.05, 0.2, 0.15, and 0.6, respectively. Missing values are imputed using the 25th percentile, and the scores are scaled to ensure that the ratio of the maximum to the minimum influence score remains within a factor of 10. All GPU-related inferences are executed on $1002\times A$ 100 GPUs.

### v4

@@ Appendix C Fact-Checking Claim Rewriting Factor
As shown in Table 9, six critical factors are proposed to systematically transform social media headlines into verifiable claims: (1) Temporal Anchoring converts vague temporal expressions into specific time references; (2) Data Granularity decomposes aggregated data into measurable units; (3) Ambiguity Resolution replaces probabilistic terms with deterministic statements; (4) Comparative Standard introduces quantifiable benchmarks; (5) Domain Knowledge integrates industry-specific parameters; and (6) Source Implication embeds provenance cues. These factors collectively enhance claim verifiability while preserving semantic coherence for automated processing. Concrete rewriting examples are illustrated in Figure 8.
@@ Appendix E Evidence Library Construction Process
First, we employ GPT-4.1 to extract multiple key-element phrases from the original claims and to generate rewritten statements based on the original claims. Figure 10 showcases the specific rewriting prompt. Specifically, we transform original statements into challenging variants by maintaining core fact relevance (relevance score >0.6) and introducing at least three types of semantic alterations: subject displacement (e.g., institution → individual), causal inversion (e.g., "A leads to B" → "B triggers A"), degree conversion (e.g., "significant growth" → "slight fluctuation"), spatiotemporal shift (across years/regions), and quantification method change (absolute value → percentage), allowing for reasonable logical leaps. The rewritten claims are then processed through Bing web retrieval, retaining the top 10 search results and extracting their textual content. Subsequently, we merge and deduplicate the newly collected evidence with pre-existing gold evidence, followed by publication date crawling for each webpage to finalize the evidence library.
Appendix F Details of the HCPI
Equation 2 provides a detailed definition of the HCPI scoring mechanism. The descriptions of the three sample types are as follows: Support (SUP): Samples originally labeled as support; Refute (REF): Samples originally labeled as refute; Not Enough Information (NEI): Samples where there is insufficient evidence to determine the authenticity of the claim.
And, the social influence score calculations for them are calculated as follows:
$S_{\text{SUP}}=\begin{aligned} &\sum_{\mathclap{i\in\text{SUP}}}\,\begin{cases}\text{s}_{i}\cdot\mathrm{ECS}(i),&\hat{y}_{i}=\text{SUP}\\ -2\cdot\text{s}_{i},&\hat{y}_{i}=\text{REF}\end{cases}\end{aligned}$ (5)
$S_{\text{REF}}=\begin{aligned} \sum_{\mathclap{j\in\text{REF}}}\,\text{s}_{j}\cdot\mathrm{ECS}(j)\cdot\delta(\hat{y}_{j}=\text{REF})\end{aligned}$ (6)
$S_{\text{NEI}}=\begin{cases}-\text{s}_{k},&\hat{y}_{k}=\text{REF}\\ \text{s}_{k}\cdot\mathrm{ECS}(k),&\hat{y}_{k}=\text{NEI}\end{cases}$ (7)
For the SUP (Support) samples, if they are correctly predicted, their social influence score is calculated as the product of Inf(i) and ECS(i). However, if the prediction is incorrect (i.e., predicted as REF), a higher penalty is applied, represented by -2 * Inf(i). This ensures that misclassifying a positive sample incurs a larger penalty. For the REF (Refute) samples, the social influence score is computed using the product of Inf(j), ECS(j), and the delta function $\delta(\hat{y}_{j}=\text{REF})$. This term captures the influence of REF samples, considering their relevance and correctness in prediction. For the NEI (Not Enough Information) samples, if misclassified as REF, the score is penalized as -Inf(k). If correctly classified as NEI, the score is calculated as Inf(k) * ECS(k), ensuring that the contribution to the score is aligned with the quality of the explanatory content.
@@ Appendix G Prompt of FactISR
Figure 11 shows the prompt of FactISR.
@@
Appendix H Example of FactISR
Figure 13 illustrates an example of FactISR. Without ISR, the model directly outputs a conclusion of insufficient evidence and prematurely ends the reasoning process. Our reflection mechanism encourages the model to reassess its previous judgment, leading to a reconsideration that ultimately results in the correct outcome.
Figure 13: An Example of FactISR.
@@ Appendix I Experiments Under Gold Evidence Conditions
We conduct experiments on PyTorch [^2] https://pytorch.org/ and $2\times A100$ GPUs. The evaluation for ESC was conducted using GPT-4.1, while BERTScore evaluations are conducted on chinese-bert-Base [^3] https://huggingface.co/google-bert/bert-base-chinese. The small constant $\epsilon$ to prevent division by zero, the hyperparameter $k$ controlling the reflection strength, and the decay factor $\gamma$ are set to $1\times 10^{-8}$, 1 and 0.5, respectively. The maximum number of reflections is set to 3. The maximum input length is set to 16k, while the maximum output lengths for LLMs and RLMs are set to 300 and 5k, respectively. The maximum length of retrieved results is truncated at 3k. The maximum number of retrievals is set to 3, and to ensure fair comparison, the maximum number of dynamically added evidences by our DEA is also limited to the same. All inference experiments utilized greedy search as the strategy. In this paper, for the HCPI metric, the values of $\alpha$, $\beta$, $\kappa$ and $\lambda$ used to calculate the influence score are set to 0.05, 0.2, 0.15, and 0.6, respectively. Missing values are imputed using the 25th percentile, and the scores are scaled to ensure that the ratio of the maximum to the minimum influence score remains within a factor of 10.